A box ha 3 netball and 6 volleyball card. What is the probability of selecting ....





A. A netball card, keeping it out, and then selecting another netball card?





B. A netball card, keeping it out, and then selecting another volleyball card?

1
Expert's answer
2022-03-17T07:22:36-0400

A. "P={\frac 3 {3+6}}*{\frac {3-1} {3+6-1}}={\frac 1 3}*{\frac 1 4}={\frac 1 {12}}"

B. "P={\frac 3 {3+6}}*{\frac 6 {3+6-1}}={\frac 1 3}*{\frac 3 4}={\frac 1 4}"